OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)). -40x + 24 -40x + 31 -40x + 64 -40x + 70

OpenStudy (anonymous):

Do you know how to start it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Alright g is on the inside, so you are going to put the function g, into the x of the function f

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does that make sense? Are you able to do it now?

OpenStudy (anonymous):

i dont even know what im looking at

OpenStudy (anonymous):

haha alright here, -4(10x-6) + 7 I put g(x) into f(x) now just solve

OpenStudy (anonymous):

simplify not solve

OpenStudy (anonymous):

how do i do that?

OpenStudy (jasmineflvs):

first analyze it the question is find f(g(x)) you already know that f(x) = -4x + 7 and g(x) = 10x - 6...but in the question there is no f(x) but there is just an f so what would you do to find f

OpenStudy (anonymous):

simplify? just distribute the parenthesis

OpenStudy (jasmineflvs):

no not yet

OpenStudy (anonymous):

yes you would, from what I said he would

OpenStudy (anonymous):

I already went through putting the function g into the function f

OpenStudy (jasmineflvs):

you would take away the x from the equation so it would be f=-4+7

OpenStudy (jasmineflvs):

i think

OpenStudy (anonymous):

;__; what do i do

OpenStudy (anonymous):

no you would not

OpenStudy (anonymous):

listen to me, you have f(x) = -4(10x-6) + 7

OpenStudy (jasmineflvs):

i'm sorry i'll try to find a link that can help you..kk..first what kind of problem is this

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

then you just simplify f(x) = -4(10x-6) + 7 -40x + 24 + 7

OpenStudy (jasmineflvs):

i mean like equation....ect.

OpenStudy (anonymous):

f(x) = -40x + 31, and that is your answer

OpenStudy (anonymous):

so b

OpenStudy (jasmineflvs):

lets have two opinions first kk

OpenStudy (anonymous):

No it is b I promise you

OpenStudy (anonymous):

im going to cry

OpenStudy (anonymous):

jasmineflvs I came up with one of the answers, you told him to eliminate the x which you DO NOT DO. I promise you it is b, I can show you again if you like

OpenStudy (jasmineflvs):

no lets get two opinions and dont cry

OpenStudy (anonymous):

Why does he need two opinions? alright thebroski I want you to work it out yourself, start by putting g(x) into the x in f(x)

OpenStudy (anonymous):

f(x) = -4x + 7 and g(x) = 10x - 6, so g(x) goes into f(x = -4(g(x) goes here) + 7

OpenStudy (jasmineflvs):

just a sec kkk i'll get it and if you were rigght ill give you a medal..kk

OpenStudy (anonymous):

thats racists

OpenStudy (jasmineflvs):

???

OpenStudy (jasmineflvs):

how?? may i ask

OpenStudy (anonymous):

.just a sec ((kkk))) i'll get it and if you were rigght ill give you a medal..kk

OpenStudy (jasmineflvs):

??? still dont understand why it is racist

OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (jasmineflvs):

let me guess you got the answer

OpenStudy (anonymous):

no im just terrible at math

OpenStudy (anonymous):

f^-1(x) is an inverse equation

OpenStudy (jasmineflvs):

that's okay

OpenStudy (anonymous):

you need to switch the spots of x and y in the equation given and the solve for y, make sense

OpenStudy (anonymous):

?

OpenStudy (jasmineflvs):

sorry i can't help you on functions i am behind in Algebra...I'm in 10th grade and I'm still in Algebra 1...sorwie

OpenStudy (anonymous):

im in 10th to but math is not my subject i dont understand any of it except 2+2 thats all i understand

OpenStudy (jasmineflvs):

lolz

OpenStudy (jasmineflvs):

i'm getting a friend to help you on this one

OpenStudy (jasmineflvs):

@sasogeek @Hero @jim_thompson5910 @cwrw238 plz help him

OpenStudy (usarmy3947):

ok got the answer

OpenStudy (usarmy3947):

here f(10x-6)=31-40x

OpenStudy (anonymous):

wut

OpenStudy (usarmy3947):

thats the answer

OpenStudy (anonymous):

oh right

OpenStudy (usarmy3947):

sorry sir if u want help be more nicer

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

.Let f(x) = 2x + 2. Solve f-1(x) when x = 4. 1 3 4 10

OpenStudy (sasogeek):

thebroski, if you're not going to put some effort into understanding the concept of the answer, no one can really help you. stop posting questions with multiple choice answers if you're not ready to learn

OpenStudy (usarmy3947):

dont say any thing without thinking this is a Moderator

OpenStudy (anonymous):

i am trying to put effort into this its just that i have a problem with understanding things im just kinda slow.

OpenStudy (anonymous):

btw i really dont care if ur the moderator it doesnt give you the right to ban someone because they need help with questions

OpenStudy (sasogeek):

i'm not going to ban u for asking for help, in fact i'll help you if i can. it just doesnt seem like u really want the help though. sounds like u just want the answer cos it's quite a long thread and u still don't get it... reposting the question with the possible answers doesn't get u so far on this site. be patient, and try at least, to understand :)

OpenStudy (anonymous):

ok

OpenStudy (usarmy3947):

ya i could also help u

OpenStudy (anonymous):

Tango Hotel Alpha November Kilo Yankee Oscar Uniform

OpenStudy (usarmy3947):

what?

OpenStudy (anonymous):

it means thank you

OpenStudy (usarmy3947):

no problem ill be leaving now bye